OpenStudy (anonymous):

Last year Elissa contributed $105 per month toward her 401(k) account. If her employer matched 15% of her contributions, what was the total amount contributed to the 401(k) at the end of the year?

OpenStudy (anonymous):

A. $1575 B. $1162 C. $1442 D. $1449

OpenStudy (anonymous):

amount she contributed = 105 * 12 = 1260 her employers contributions per month = 15/100 * 105 = 15.75 amount employer contributed for the whole year = 15.75* 12 = 189 total = 1260 + 189 = 1449 (D)